OpenStudy (anonymous):

Look at this linear equation. 3x + 5y = 30 Which of the following equations can be used to solve for y in the linear equation above?

OpenStudy (anonymous):

what are the following equations?

OpenStudy (anonymous):

A.y=-3/5x+30 B.y=3x+6 C.y=-3/5x=6 D.=3/5x+30

OpenStudy (anonymous):

y=(-3/5)x+6

OpenStudy (anonymous):

C is the answer

OpenStudy (anonymous):

C.y=-3/5x=6 3x + 5y = 30 ->y = (-3/5)x + 30/5 = (-3/5)x + 6